OpenStudy (littlebird):

How would I algebraicly find the x intercept of y=(2^-x)-1?

OpenStudy (anonymous):

x-intercepts are points of the form \((x,0)\). So for this value of x, the value of y is 0. So you solve the following equation: \[0=2^{-x}-1\]

OpenStudy (littlebird):

I don't understad what to do after I get \[2^{-x} = 1\]

OpenStudy (anonymous):

\[2^{-x}=\frac{1}{2^x}=1\] which is the same as \[1=2^x\] easier?